Derek was 12 years old when his sister was born. Now the sum of their ages is 58. How old is Derek now?
Answer by erica65404(394)

D=12+s
s+D=58
plug in D
s+12+s=58
2s+12=58
2s=46
s=23
plug into D=12+s
D=12+23
D=35
Check answer
35+23=58
Derek is 35 and his sister is 23